OpenStudy (anonymous):

Solve the equation. Check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

please help me do this problem

OpenStudy (anonymous):

well first subtract 3 from both sides

OpenStudy (anonymous):

so -2/x+1 = 4/x?

OpenStudy (anonymous):

now add 2/x on both sides

OpenStudy (anonymous):

4 = 6/x+3?

OpenStudy (anonymous):

hold on is your original equation this \[\frac{ 4 }{x+3 }\]

OpenStudy (anonymous):

or is it the way you have it?

OpenStudy (anonymous):

it's the way I had it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

any way how did you get 4=6/x+3

OpenStudy (anonymous):

wasn't I supposed to add 2/x on both sides?

OpenStudy (anonymous):

yes but what is -2/x+2/x

OpenStudy (anonymous):

where did that come from?

OpenStudy (anonymous):

ok so let's go back we had \[\frac{ -2 }{ x }+1=\frac{ 4 }{ x }\] if we add 2/x to both sides it looks like this\[\frac{ -2 }{ x }+1+\frac{ 2 }{ x }=\frac{ 4 }{ x }+\frac{ 2 }{ x }\]

OpenStudy (anonymous):

now combine like terms

OpenStudy (anonymous):

actually the problem is different, let me make the equation

OpenStudy (anonymous):

\[-2/x+4 = 4/x+3\]

OpenStudy (anonymous):

is it \[\frac{ -2 }{ x+4 }=\frac{ 4 }{ x+3 }\]

OpenStudy (anonymous):

yes, I'm not sure how to exactly make it like that though lol

OpenStudy (anonymous):

just to make it more clear next time just put everything that is in the bottom portion of the fraction in parentheses like this -2/(x+4)

OpenStudy (anonymous):

but any back to the problem. Now you want to cross multiply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me know what you got after cross multiplying

OpenStudy (anonymous):

(-2x+3)/4x+4?

OpenStudy (anonymous):

that's close but you forgot to multiply 4*4 on the right side and 3*3 on the left side also they should equal each other not be in a fraction. \[3(x+3)=4(x+4)\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

now distribute the 3 on the left side and the 4 on the right

OpenStudy (anonymous):

would we subtract 3 from both sides first?

OpenStudy (anonymous):

no just distribute each side first

OpenStudy (anonymous):

Do you remember the distributive property

OpenStudy (anonymous):

no, how was it again?

OpenStudy (anonymous):

a(b+c)=a*b+a*c

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3*x+3*3 = 4*x+4*4?

OpenStudy (anonymous):

ok just simplify that. What is 3*3 and what is 4*4

OpenStudy (anonymous):

9 and 16

OpenStudy (anonymous):

3*x+9 = 4*x+16

OpenStudy (anonymous):

now subtract 3x from both sides

OpenStudy (anonymous):

9 = x+16

OpenStudy (anonymous):

what do you think comes next

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

but it doesn't match the original equation

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

oh wait -7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you mean it doesn't match the original equation

OpenStudy (anonymous):

if you substitute -7 for x, it won't equal the same amount for both sides

OpenStudy (anonymous):

do you know why?

OpenStudy (anonymous):

oh because earlier i accidentally wrote in 3(x+3) instead of -2(x+3)

OpenStudy (anonymous):

so just solve for -2(x+3)=4(x+4)

OpenStudy (anonymous):

distributive property?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2*x+-2*3 = 4*x+4*4?

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

add 2x to both sides?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh wait, simplify

OpenStudy (anonymous):

yea simplify

OpenStudy (anonymous):

-2x -6 = 4x+16

OpenStudy (anonymous):

then -6 = 6x+16?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then add 6 to both sides?

OpenStudy (anonymous):

or subtract -16 from both sides?

OpenStudy (anonymous):

-22=6x?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so -11/3

OpenStudy (anonymous):

yes and that is definitely correct because I checked

OpenStudy (anonymous):

ok, thank you...I'll give you best response =)

OpenStudy (anonymous):

LMAO you better I'm the only responder

OpenStudy (anonymous):

lol, could you help me with one last word problem?

OpenStudy (anonymous):

ok but we're gonna do this one faster because an hour was way too long for the last problem

OpenStudy (anonymous):

yeah it should be much quicker

OpenStudy (anonymous):

The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles each way against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs?

OpenStudy (anonymous):

What is this a physics problem?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

dude I'm not to sure about this one sorry.

OpenStudy (anonymous):

oh, that's how I felt when I read it =/

OpenStudy (anonymous):

Good Luck

OpenStudy (anonymous):

thanks, I'll try to open a question with it and have a great day :)

OpenStudy (anonymous):

you 2